OpenStudy (miguel980):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (anonymous):

yes

OpenStudy (miguel980):

I need an example also and step by step instructions.

OpenStudy (anonymous):

using tangent = opp/adj, cosine = adj/hyp and sine = opp/hyp

OpenStudy (anonymous):

are you allowed to use calculators?

OpenStudy (anonymous):

that would give the angle of 1 side. and since the other side is 90 degrees, you can then find the last angle

OpenStudy (miguel980):

Yes I can use a calculator.

OpenStudy (anonymous):

eg

OpenStudy (miguel980):

So either formula could be used?

OpenStudy (anonymous):

just follow irkiz

OpenStudy (anonymous):

any formula can be used depending which side you have the length

OpenStudy (anonymous):

need an example?

OpenStudy (miguel980):

Sure

OpenStudy (anonymous):